# Beauville-Bogomolov decomposition for klt varieties

### v1

@@
Henri Guenancia Institut de Mathématiques de Toulouse; UMR 5219, Université de Toulouse; CNRS, UPS, 118 route de Narbonne, F-31062 Toulouse Cedex 9, France henri.guenancia@math.cnrs.fr
@@ 1.4. Brief outline of the proof
Acknowledgements. These lecture notes are based on a series of lectures given at the C.I.M.E summer school on Calabi-Yau varieties in 2024. I am very grateful to the organizers Simone Diverio, Vincent Guedj, and Hoang Chinh Lu for the kind invitation to participate. I am indebted to Hyunsuk Kim and the anonymous referee for their very thorough and careful reading of these notes and for their numerous valuable remarks and suggestions which helped improve the exposition. Finally, I would like to thank my collaborators Ben Bakker, Junyan Cao, Benoît Claudon, Stéphane Druel, Patrick Graf, Daniel Greb, Stefan Kebekus, Christian Lehn and Mihai Păun for all the fruitful discussions we have had over the years on the topics appearing in these notes.
@@ 6. Irreducible pieces of the decomposition
This example tells us that the geometry (or, roughtly speaking the type) of a given variety can drastically change when one passes to a quasi-étale, non-étale cover. This suggests that whether one looks at $\pi_{1}(X)$ or $\pi_{1}(X_{\rm reg})$ will provide vastly distinct geometric insight into $X$. Many more relevant examples of singular varieties with trivial canonical bundle have been considered in the past; a thorough account was provided in [GGK19, Section 14] with a view towards classification.
@@ 9.1. Fundamental groups
Note that Conjecture 9.1 is a consequence of the more general abelianity conjecture formulated by Campana).
@@ 9.3. Varieties with nef anticanonical bundle
In the last few decades, the classification of projective or Kähler manifolds with nef anticanonical class has drawn a lot of attention, starting at least with seminal work by Demailly-Peternell-Schneider [DPS94, DPS]. A major result was proved in the projective case by Cao-Höring [CH19] relying on earlier progress by Cao [Cao19].
Theorem 9.6 ([CH19]).
Let $X$ be a projective manifold such that $K_{X}$ is nef. Then, there exists a locally trivial fibration
@@
One expects that the result holds for Kähler manifolds as well but that case is still open in general. It was recently solved for Kähler threefolds by Matsumura-Wu [MW24].
Given the Beauville-Bogomolov decomposition theorem, Theorem 9.6 implies that if $X$ is projective with $K_{X}$ nef, then the universal cover $\widetilde{X}$ of $X$ is isomorphic to $F\times{\mathbb{C}}^{r}\times\prod_{i\in I}Y_{i}\times\prod_{j\in J}Z_{j}$ where $Y_{i}$ are ICY manifolds and $Z_{j}$ are IHS manifolds.
The structure theorem above was generalized to projective klt pairs $(X,\Delta)$ with $(K_{X}+\Delta)$ nef by Matsumura-Wang [MW23] after important work by Campana-Cao-Matsumura [CCM21]. The result reads as follows:
@@ 9.4. No Beauville-Bogomolov decomposition for log canonical singularities
  • [DPS94] J.-P. Demailly, T. Peternell & M. Schneider – « Compact complex manifolds with numerically effective tangent bundles », J. Algebr. Geom. 3 (1994), no. 2, p. 295–345 (English).
@@
  • [Dru18] S. Druel – « A decomposition theorem for singular spaces with trivial canonical class of dimension at most five », Invent. Math. 211 (2018), no. 1, p. 245–296.
@@
  • [GP24] H. Guenancia & M. Păun – «  Bogomolov-Gieseker inequality for log terminal Kähler threefolds  », Preprint arXiv:2405.10003, 2024.
@@
  • [HP19] A. Höring & T. Peternell – « Algebraic integrability of foliations with numerically trivial canonical bundle », Invent. Math. 216 (2019), no. 2, p. 395–419.
@@
  • [MW23] S.-i. Matsumura & J. Wang – « Structure theorem for projective klt pairs with nef anti-canonical divisor », arXiv:2105.14308, 2023.
  • [MW24] S.-i. Matsumura & X. Wu – « Compact Kähler 3-folds with nef anti-canonical bundle », arXiv:2304.03163, to appear in Math. Ann., 2024.

### v2

@@
Henri Guenancia Univ. Bordeaux, CNRS, Bordeaux INP, IMB, UMR 5251, F-33400 Talence, France henri.guenancia@math.cnrs.fr
@@ 1.4. Brief outline of the proof
Acknowledgements. These lecture notes are based on a series of lectures given at the C.I.M.E summer school on Calabi-Yau varieties in 2024. I am very grateful to the organizers Simone Diverio, Vincent Guedj, and Hoang Chinh Lu for the kind invitation to participate. I am indebted to Hyunsuk Kim and the anonymous referees for their very thorough and careful reading of these notes and for their numerous valuable remarks and suggestions which helped improve the exposition. Finally, I would like to thank my collaborators Ben Bakker, Junyan Cao, Benoît Claudon, Stéphane Druel, Patrick Graf, Daniel Greb, Stefan Kebekus, Christian Lehn and Mihai Păun for all the fruitful discussions we have had over the years on the topics appearing in these notes.
@@ 6. Irreducible pieces of the decomposition
This example tells us that the geometry (or, roughly speaking the type) of a given variety can drastically change when one passes to a quasi-étale, non-étale cover. This suggests that whether one looks at $\pi_{1}(X)$ or $\pi_{1}(X_{\rm reg})$ will provide vastly distinct geometric insight into $X$. Many more relevant examples of singular varieties with trivial canonical bundle have been considered in the past ; a thorough account was provided in [GGK19, Section 14] with a view towards classification.
@@ 2. $\bullet$
In the past, many different definitions of what a singular holomorphic symplectic variety could be were proposed, cf [GGK19, Section 14] for a discussion on that topic, or Section 9.2.1 below for the connection with the notion of primitive symplectic variety. More recently, [GPP24] discussed the surface case at length while [BGMM25] provided a whole host of old and new examples coming out of performing terminalization of quotients of IHS manifolds. Finally, let us point out to [PR18] and [Sac23] who give interesting examples of IHS varieties not admitting a symplectic resolution ; these example arise from the study of holomorphic moduli spaces of sheaves and Bridgeland-stable objects.
@@ 9.1. Fundamental groups
Note that Conjecture 9.1 is a consequence of the more general abelianity conjecture formulated by Campana.
@@ 9.3. Varieties with nef anticanonical bundle
In the last few decades, the classification of projective or Kähler manifolds with nef anticanonical class has drawn a lot of attention, starting at least with seminal work by Demailly-Peternell-Schneider [DPS94, DPS]. The following major result was first proved for projective manifolds by Cao-Höring [CH19] relying on earlier progress by Cao [Cao19]. The general Kähler case was later established by Matsumura-Wang-Wu-Zhang [MWWZ25]. More precisely, they followed the general strategy of Cao-Höring and capitalized on the breakthrough [Ou25] (see also [CP25]) combined with the results in [CH24].
Theorem 9.6 ([CH19, MWWZ25]).
Let $X$ be a compact Kähler manifold such that $K_{X}$ is nef. Then, there exists a locally trivial fibration
@@
Given the Beauville-Bogomolov decomposition theorem, Theorem 9.6 implies that if $X$ is compact Kähler with $K_{X}$ nef, then the universal cover $\widetilde{X}$ of $X$ is isomorphic to $F\times{\mathbb{C}}^{r}\times\prod_{i\in I}Y_{i}\times\prod_{j\in J}Z_{j}$ where $Y_{i}$ are ICY manifolds and $Z_{j}$ are IHS manifolds.
The above structure theorem was extended to projective klt pairs $(X,\Delta)$ with $(K_{X}+\Delta)$ nef by Matsumura-Wang [MW23] after important work by Campana-Cao-Matsumura [CCM21]. The result reads as follows :
@@ 9.4. No Beauville-Bogomolov decomposition for log canonical singularities
  • [BGMM25] V. Bertini, A. Grossi, M. Mauri & E. Mazzon – « Terminalizations of quotients of compact hyperkähler manifolds by induced symplectic automorphisms », Épijournal de Géom. Algébr., EPIGA 9 (2025), p. 53 (English), Id/No 14.
@@
  • [CH24] B. Claudon & A. Höring – « Projectivity criteria for Kähler morphisms », Preprint arXiv :2404.13927, 2024.
@@
  • [CP25] J. Cao & M. Păun – « Remarks on relative canonical bundles and algebraicity criteria for foliations in Kähler context », Preprint arXiv :2502.02183, 2025.
@@
  • [DPS94] by same author, « Compact complex manifolds with numerically effective tangent bundles », J. Algebr. Geom. 3 (1994), no. 2, p. 295–345 (English).
@@
  • [Dru18] by same author, « A decomposition theorem for singular spaces with trivial canonical class of dimension at most five », Invent. Math. 211 (2018), no. 1, p. 245–296.
@@
  • [GP24] H. Guenancia & M. Păun – «  Bogomolov-Gieseker inequality for log terminal Kähler threefolds », Preprint arXiv :2405.10003, to appear in Comm. Pure Appl. Math., 2024.
  • [GPP24] A. Garbagnati, M. Penegini & A. Perego – « Singular symplectic surfaces », arXiv :2407.21173, 2024.
@@
  • [HP19] by same author, « Algebraic integrability of foliations with numerically trivial canonical bundle », Invent. Math. 216 (2019), no. 2, p. 395–419.
@@
  • [MW23] S.-i. Matsumura & J. Wang – « Structure theorem for projective klt pairs with nef anti-canonical divisor », arXiv :2105.14308, to appear in JEMS, 2023.
  • [MWWZ25] S.-i. Matsumura, J. Wang, X. Wu & Q. Zhang – « Compact Kähler manifolds with nef anti-canonical bundle », arXiv :2506.23218, 2025.
@@
  • [Ou25] W. Ou – « A characterization of uniruled compact Kähler manifolds », Preprint arXiv :2501.18088, 2025.
@@
  • [PR18] A. Perego & A. Rapagnetta – « The moduli spaces of semistable sheaves on K3 surfaces are irreducible symplectic varieties », Preprint arXiv :1802.01182, to appear in Algebraic Geometry, 2018.
@@
  • [Sac23] G. Saccà – « Singular symplectic surfaces », arXiv :2304.02609, to appear in "Current developments in Hodge theory. Proceedings of Hodge theory at IMSA”, Simons Symposia Series, 2023.